If 500g of marble are decomposed completely and knowing that it contains 90% calcium carbonate, calculate the mass of calcium oxide and the volume of carbon dioxide ( at S.T.P.) produced.
1
Expert's answer
2020-05-19T08:59:05-0400

The molar mass of CaCO3 is 100 g/mol

The molar mass of CaO is 56 g/mol

The molar volume of a gas at STP is 22.4 L/mol


Solution:

m(CaCO3) = m(marble) × w = 500 g × 0.9 = 450 g


The balanced chemical equation:

CaCO3 --> CaO + CO2

According to the chemical equation: n(CaCO3) = n(CaO) = n(CO2)


Moles of CaCO3 = n(CaCO3) = Mass of CaCO3 / Molar mass of CaCO3

n(CaCO3) = (450 g) / (100 g/mol) = 4.5 mol

n(CaCO3) = n(CaO) = n(CO2) = 4.5 mol


Mass of CaO = Moles of CaO × Molar mass of CaO

m(CaO) = (4.5 mol) × (56 g/mol) = 252 g

m(CaO) = 252 g


Volume of CO2 = Moles of CO2 × Molar volume of CO2

V(CO2) = (4.5 mol) × (22.4 L/mol) = 100.8 L

V(CO2) = 100.8 L


Answer:

The mass of calcium oxide (CaO) is 252 g;

The volume of carbon dioxide (CO2) is 100.8 L.